OpenStudy (roxy.girl):

2a + 5 = 9a + -16? solve

OpenStudy (roxy.girl):

@Qwertty123. @Qwertty123 so you would add -9?

OpenStudy (qwertty123.):

yes. :)

OpenStudy (qwertty123):

yes. Add '-9a' to each side of the equation. 5 + 2a + -9a = -16 + 9a + -9a Combine like terms: 2a + -9a = -7a 5 + -7a = -16 + 9a + -9a

OpenStudy (qwertty123.):

So...Combine like terms: 9a + -9a = 0 5 + -7a = -16 + 0 5 + -7a = -16

OpenStudy (roxy.girl):

okay

OpenStudy (roxy.girl):

so you add negative 5?

OpenStudy (qwertty123):

yes!

OpenStudy (qwertty123.):

yes

OpenStudy (roxy.girl):

so 5 + -5 + -7a = -16 + -5

OpenStudy (qwertty123):

Combine like terms: 5 + -5 = 0 0 + -7a = -16 + -5 -7a = -16 + -5 Combine like terms: -16 + -5 = -21 -7a = -21

OpenStudy (roxy.girl):

Qwertty you divide by -7 right?

OpenStudy (qwertty123.):

yes

OpenStudy (qwertty123):

so you get?

OpenStudy (roxy.girl):

3!

OpenStudy (roxy.girl):

thanks girls

OpenStudy (qwertty123.):

yes!

OpenStudy (qwertty123):

you welcome